OpenStudy (anonymous):

Let F(x) =x^4e^x. Determine the nth derivative of f at x=0

OpenStudy (turingtest):

taylor series?

OpenStudy (anonymous):

no

OpenStudy (turingtest):

\[f(x)=x^4e^x\]right?

OpenStudy (anonymous):

\(\huge F(x) =x^4e^x\) or \(\huge F(x) =x^{4e^x}\) ???

OpenStudy (anonymous):

Yes turingtest that is correct

OpenStudy (turingtest):

really? the first one that agentx5 has?

OpenStudy (turingtest):

because the second makes a lot more sense

OpenStudy (anonymous):

Product rule! eh? :-)

OpenStudy (anonymous):

It is the first one that agentx5 has

OpenStudy (anonymous):

First d Second + Second d First ^_^ (this is the way I say it)

OpenStudy (anonymous):

hmm just for the first part \[ f'(x)=4x^3e^x+e^xx^4 \] If we derive more and more the first term will become \[ n! e^x \] right?

OpenStudy (anonymous):

\[\large \frac{d}{dx}(x^4)(e^x) = (x^4)\frac{d}{dx}(e^x) + (e^x)\frac{d}{dx}(x^4)\] \[\large f'(x) = (x^4)(e^x) + (e^x)(4x^3)\]

OpenStudy (turingtest):

\[f'(x)=4x^3e^x+x^4e^x\]\[f''(x)=12x^2e^x+4x^3e^x+4x^3e^x+x^4e^x=12x^2e^x+8x^3e^x+x^4e^x\]\[f'''(x)=24xe^x+12x^2e^x+24x^2e^x+8x^3e^x+4x^3e^x+x^4e^x\]\[f^{(4)}(x)=24e^x+O(x)\]so only from the 4th derivative on do we get a non zero value for the nth derivative of f at 0

OpenStudy (turingtest):

I wonder if we can apply\[(uv)'=\sum_{k=0}^n\binom nku^{(n-k)}v^{(k)}\]to get a more general formula...

OpenStudy (anonymous):

You can write the first derivative also as: \(e^x x^3 (x+4)\) In which case the second is... \(e^x x^2 (x^2+8 x+12)\) Third is... \(e^x x^2 (x^2+8 x+12)\) Fourth is... \(e^x x (x^3+12 x^2+36 x+24)\) Fifth is... \(e^x (x^4+16 x^3+72 x^2+96 x+24)\) Sixth is... \(e^x (x^4+20 x^3+120 x^2+240 x+120)\) etc...

OpenStudy (anonymous):

Oh for crying out loud... I wrote the second and third twice... err

OpenStudy (anonymous):

But yeah @TuringTest , chain rule and product rule says your polynomial stuff doesn't go away... Right?

OpenStudy (anonymous):

You keep cycling an x back into it, almost like recursion

OpenStudy (turingtest):

go away? not sure what you mean... you mean the O(x) part yes it's only true for the fourth derivative, for the others it will come back into play

OpenStudy (anonymous):

Yeah what's with the O(x) ?

OpenStudy (anonymous):

It doesn't go to zero, does it?

OpenStudy (anonymous):

Wait '0' or 'O'?

OpenStudy (turingtest):

that means that they are "junk" terms i.e. they vanish when you plug in x=0 because they all are being multiplied by x

OpenStudy (anonymous):

Junk terms? :-D I wonder if I can do that on an exam :-3

OpenStudy (turingtest):

but I didn't mean to imply that they can be ignored from that point on; they cannot and will need to be accounted for int the higher derivatives

OpenStudy (anonymous):

"Sry, this section of the test is junk" "Divergent by the Turing Junk Theorem" :-3

OpenStudy (turingtest):

sure that notation is done a lot in calc proofs using the difference quotient (proof of the power rule for example involves a bunch of junk terms that have an h in them, which dissappear as \(h\to0\) and are usually written as \(O(x)\) to save paper)

OpenStudy (turingtest):

haha not quite that I don't think :P I have a good like where an MIT prof uses it of I can find it...

OpenStudy (anonymous):

Hmm after 4 derivatives the first term becomes \(n!e^x\) therefore a number, at x=0, after 5 derivates the first two terms become \(n!e^x\) and therefore a number for x=0, or are my derivatives wrong?

OpenStudy (anonymous):

but yeh, it depends on more then just n.

OpenStudy (experimentx):

i guess we use the method Turing suggested.

OpenStudy (turingtest):

@agentx5 check out around 47:00 on here in the proof of the power rule to see how it is used properly http://ocw.mit.edu/courses/mathematics/18-01sc-single-variable-calculus-fall-2010/part-a-definition-and-basic-rules/session-2-examples-of-derivatives/

OpenStudy (turingtest):

...the Junk think I mean @experimentX which method? the leibniz rule that I wrote up there? (I think that's what it's called at least....)

OpenStudy (anonymous):

Err which one of the 4? (also: yikes!)

OpenStudy (experimentx):

yes ...it's pretty useless to expand.

OpenStudy (experimentx):

 all those stupid 24's keep piling up. it's better this way.